A uniform metallic chain in form of circular loop of mass m with a length l rotates at the rate of n revolution per sec Find tension T in the chain

Grade:11

1 Answers

Rahul
6 years ago
angular speed, w=n*2pi rad/snow, tension = centripetal force therefore mw^2r=tension(T) hence, ml(2npi)=T (r=l)